Given f(x)=2x^2-3x+5, Find f(1).
Answer by checkley71(8403)

F(X)=2X^2-3X+5
F(1)=2*1^2-3*1+5
F(1)=2*1-3*1+5
F(1)=2-3+5
F(1)=4 ANSWER.